A farmer wishes to fence his rectangular paddock, if its measurements are 366 metres long and 100 metres wide, what is the perimeter?
Answer by satyareddy22(84)

length = 366m, breadth = 100m
perimeter=2(l+b)=2(366+100)=2(466)=932m